weoutherebored:

-13x= -195 fined the equation

mxddi3:

Welcome to QC!! Are you asking how to solve this bc i'm not sure what you mean by "fined the equation"

weoutherebored:

solve the problem

mxddi3:

Oh, alright. This is quite simple. We must isolate the variable and get it on the left side by itself. To do so, we divide both sides by -13 (:

weoutherebored:

k thanks

boredfr:

Step 1: Divide both sides by -13. −13x −13 = −195 −13 x=15

weoutherebored:

k